OpenStudy (anonymous):

A spherical balloon with radius r inches has volume V(r) = 4 3 πr3. Find an expression that represents the amount of air required to inflate the balloon from a radius of r inches to a radius of r + 1 inches. (Express your answer in terms of π and r.)

OpenStudy (anonymous):

\[ V(r)= \frac 43 \pi r^3\\ V(r+1)= \frac 43 \pi(r+1)^3\\ V(r+1) - V(r)= \frac 43 \pi(r+1)^3-\frac 43 \pi r^3=?? \]

OpenStudy (anonymous):

After expanding and simplifying, one gets \[ V(r+1)-V(r)=\frac{4}{3} \pi \left(3 r^2+3 r+1\right) \]

OpenStudy (anonymous):

In the above, we used \[ (r+1)^3=r^3+3 r^2+3 r+1 \]